THE SUBMILLIMETER SPECTRUM OF THE GROUND HYDROXYL TORSIONAL STATE OF ETHYL ALCOHOL (CH_{3}CH_{2}OH)

\underline{J. C. PEARSON}, Jet Propulsion Laboratory, California Institute of Technology, Mail Stop 183-301, 4800 Oak Grove Dr., Pasadena, CA 91109; K. V. L. N. SASTRY, Department of Physics, University of New Brunswick, Fredericton, New Brunswick, E3B 5A3 Canada; E. HERBST and F. C. DE LUCIA, Department of Physics, The Ohio State University, 174 W. 18th Ave., Columbus, OH 43210.

The ground state spectrum of the three torsional sub-states (trans, guanche+ and gauche-) of the asymmetric internal hydroxyl rotation in ethyl alcohol has been recorded through 652 GHz. Our previous trans^{a} and gauche^{b} work has been expanded to include approximately 3000 assigned transitions through J=40 and K_{a}=17 including approximately 1000 which show significant deviation from a standard asymmetric rotor pattern due to transgauche interactions. The location of these interactions has enabled us to determine the trans-gauche energy difference accurately. The application of a fixed-frame-axis method (FFAM) Hamiltonian to this asymmetric top-asymmetric frame internal rotation problem is discussed. The results of a three-state FFAM analysis are presented.

aJ. C. Pearson, K. V. L. N. Sastry, M. Winnewisser, E. Herbst and F. C. De Lucia, J. Phys. Chem. Ref. Data \underline{24}, 1 (1995).

bJ. C. Pearson, K. V. L. N. Sastry, E. Herbst and F. C. De Lucia, J. Mol. Spectrosc., in press (1996).